OpenStudy (tyteen4a03):

Given two vectors \(\vec{a} = (2, 3), \vec{b} = (4, -1)\). Project vector B onto vector A, get the new vector's coordinates and its length. I managed to find the length first but it didn't quite match up with the answer with the one in the book. How do you do these kind of questions?

ganeshie8 (ganeshie8):

projection of \(\vec{b}\) along \(\vec{a}\) = \(\large \vec{b} . \hat{a}\)

OpenStudy (tyteen4a03):

@ganeshie8 That was surprisingly easy - no idea why the resources I found overcomplicates it. Thanks!

ganeshie8 (ganeshie8):

:) yup its easy once u see it !

OpenStudy (tyteen4a03):

@ganeshie8 But that's the scalar product though. What would \(\vec{b_a}\) be then?

ganeshie8 (ganeshie8):

multiply the magnitude \(\vec{b} . \hat{a}\) wid the unit vector \(\hat{a}\)

ganeshie8 (ganeshie8):

\((\vec{b} . \hat{a} ) \hat{a}\)

ganeshie8 (ganeshie8):

that doesnt make sense is it

OpenStudy (tyteen4a03):

\(\hat{a} = (-y, x)\) according to my useless textbook. It apparently rotates the vector by 90 degrees clockwise.

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

suppose those are the two given vectors \(\vec{a}\) and \(\vec{b}\)

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

clearly vector projection of a along b = \((a\cos \theta )\hat{b} = (\vec{a} . \hat{b})\hat{b} \)

ganeshie8 (ganeshie8):

ur textbook thing makes no sense to me lol

OpenStudy (tyteen4a03):

@ganeshie8 What should \hat{a} stand for? Unit vector?

ganeshie8 (ganeshie8):

Yes ! \(\hat{a}\) is the unit vector along \(\vec{a}\)

ganeshie8 (ganeshie8):

\(\hat{b}\) is the unit vector along \(\vec{b}\)

ganeshie8 (ganeshie8):

projection formula above follows from the fact that mag of unit vector is 1....

ganeshie8 (ganeshie8):

lets work the present problem may be

ganeshie8 (ganeshie8):

\(\vec{a} = (2, 3), \vec{b} = (4, -1) \)

OpenStudy (tyteen4a03):

\[\vec{e_a}\] \[\begin{Bmatrix} \frac{2}{\sqrt{2^2 + 3^2}} \\ \frac{3}{\sqrt{2^2 + 3^2}} \end{Bmatrix}\] \[=\begin{Bmatrix} \frac{2}{\sqrt{13}} \\ \frac{3}{\sqrt{13}} \end{Bmatrix}\] \[|\vec{b_a}| = \frac{|\vec{a} \cdot \vec{b}|}{|\vec{a}|}\] \[=\frac{|(2)(4) - 3|}{\sqrt{4^2 + (-1)^2}}\] \[=\frac{5}{\sqrt{17}}\] This is what I have. e_a = unit vector of a.

ganeshie8 (ganeshie8):

you made a mistake at 2nd step from last

ganeshie8 (ganeshie8):

bottom u shud take mag of a instead, u took mag of b

ganeshie8 (ganeshie8):

\(\large |\vec{b_a}| = \frac{|\vec{a} \cdot \vec{b}|}{|\vec{a}|}\) \(\large |\vec{b_a}| = \frac{|(2)(4)-3|}{\sqrt{2^2+3^2}}\)

OpenStudy (tyteen4a03):

That was a silly mistake... Thanks :P

ganeshie8 (ganeshie8):

np :)